# Using Uncertainty Quantification to Characterize and Improve Out-of-Domain Learning for PDEs

### v1

@@ 1 Introduction
NOs have been shown to approximate the ground truth solution operator well for in-domain tasks  (Lu et al., 2019; Li et al., 2020a; Saad et al., 2023; Alesiani et al., 2022; Benitez et al., 2023). However, to be useful, they also need to be robust for out-of-domain (OOD) applications. The robustness of OOD predictions is necessary in practical applications, as the test-time PDE parameters are typically not known during training. Figure 1 shows the sharp moving (discontinuous) shock solution to the “hard” Stefan problem, where the PDE parameter $u^{*}$ denotes the solution value at the shock point. We see that while a NO model trained to map the input PDE parameter $u^{*}$ to the solution of the Stefan problem accurately captures the sharp dynamics for in-domain values of $u^{*}$, it fails for OOD values of $u^{*}$, i.e., those values of $u^{*}$ that are outside the range covered by the training data. This failure shows that NOs are not robust in OOD scenarios and thus may not effectively capture the “true” underlying operator map in practical applications. It is important when developing these models to use reliable mechanisms to detect and correct such inaccurate predictions, as this limits their widespread deployment in critical scientific applications.
@@
(b) OOD
Figure 1: 1-d Stefan Equation. In-domain (a) and OOD (b) predictions from FNO with post-hoc uncertainty estimates (3 standard deviations) obtained by BayesianNO (Magnani et al., 2022) at time $t=0.5$. The model outputs an accurate solution with reasonable uncertainty bounds in-domain (a), where the training and test input parameters are drawn from the same distribution, i.e., $u{{}^{*}}^{{\text{train}}},u{{}^{*}}^{{\text{test}}}\in[0.6,0.7]$. The model fails to do so when the test inputs are OOD (b), with $u{{}^{*}}^{{\text{test}}}\in[0.5,0.55]$. The large oscillation occurs at different locations for different seeds and we improve it.
@@ 2.2 Sensitivity Analysis for Operator Learning
$p(u^{(i)}|{\mathbf{\theta}},\phi^{(i)})=\mathcal{N}(u^{(i)};{A}(\phi^{(i)};{\mathbf{\theta}}),\frac{1}{\alpha^{2}}\|{{A}^{\dagger}}(\phi^{(i)})\|^{2}_{2}{\mathbf{I}}).\vspace{-0.2cm}$
@@
${p({\mathbf{\theta}}|{\mathcal{D}}^{\text{train}})}\propto\prod_{i=1}^{N}p(u^{(i)}|{\mathbf{\theta}},\phi^{(i)})p({\mathbf{\theta}}).\vspace{-0.2cm}$ (2)
@@ 3.2 How do Ensembles Provide Better OOD UQ
The intermediate outputs from the different models in the ensemble also show diversity in operator learning until the last layer with a key difference between the in-domain and OOD final outputs. Given the same in-domain input, the first feed-forward (lifting) layer of these models already produces diverse outputs which continues until the penultimate layer (Figure 2(c)). Figure 2(d) shows that the last layer removes the diversity and maps the outputs to the same ground truth for an in-domain input. Figures 2(e) and 2(f) illustrates that a different trend occurs for an OOD input. The OOD intermediate outputs are also diverse, the difference here is that this diversity does not vanish from the last layer outputs since the model is not able to match the ground truth data. This representative example demonstrates how the ensemble model captures epistemic uncertainty in the OOD setting. We see that the individual models in the ensemble fit the training data with close to zero error, but are diverse and disagree for OOD inputs. Ensembles can approximate the BMA in Equation 3 with functionally diverse models from the posterior, leading to better OOD uncertainty estimates.
@@ 4.1 DiverseNO: Computationally Efficient OOD UQ
For the architecture change, we modify the last feed-forward layer to have $M$ output/prediction heads instead of one. The architecture may be viewed as an ensemble with the individual models sharing parameters up to the penultimate layer. While parameter-sharing significantly reduces the computational complexity, when compared to using a full ensemble, it simultaneously hinders the diversity within the ensemble, which we showed is a crucial element for generating good uncertainty estimates. (See Section 3.2.)
To mitigate this, and to encourage diverse predictions, we propose to maximize the following diversity measure among the last-layer weights corresponding to the $M$ prediction heads. We also constrain the in-domain predictions from each of these heads to match the ground truth outputs. Formally, we solve:
@@
where ${\mathbf{\theta}}_{m}$ and ${\mathbf{\theta}}_{k}$ denote the last-layer weights corresponding to $m$-th and $k$-th prediction heads, respectively, $\hat{u}_{m}^{(i)}$ denotes the prediction from $m$-th output head for the $i$-th training example, and $u^{(i)}$ denotes the corresponding ground truth. The first term in Equation 4 is the relative $L_{2}$ loss standard in NO training. We add the regularizing second term to encourage diversity in the last-layer weights corresponding to the different head. The hyperparameter $\lambda_{\text{diverse}}$ controls the strength of the diversity regularization relative to the prediction loss. A naive selection procedure for $\lambda_{\text{diverse}}$ using only in-domain validation MSE selects an unconstrained model with no diversity penalty. Instead, we select the maximum regularization strength $\lambda_{\text{diverse}}$ that achieves an in-domain validation MSE within 10% of the identified best in-domain validation MSE. 10% is an arbitrary tolerance that denotes the % accuracy the practitioner is willing to forego to achieve better OOD UQ. This procedure trades off in-domain prediction errors for higher diversity that is useful for OOD UQ. See Appendix C for ablations with regularizations for ensemble NNs that directly diversify the predictions (Bourel et al., 2020; Zhang et al., 2020).
@@ 4.2 Operator-ProbConserv
    Use the update rule from ProbConserv, described in Section 4.2 below, to improve to the model.
@@ 5 Empirical Results
We use the Fourier Neural Operator (FNO) (Li et al., 2022a) as a base model and leave similar investigations of other operators to future work. To provide uncertainty estimates in PDE applications, we compare our proposed UQ method, DiverseNO,  [^2] The code is available at https://github.com/amazon-science/operator-probconserv. with the following four commonly-used baselines of approximate Bayesian inference (Psaros et al., 2023): (i) BayesianNO (Magnani et al., 2022), that uses the (last-layer) Laplace approximation over the MAP estimate to approximate the posterior; (ii) VarianceNO (Lakshminarayanan et al., 2017), which outputs the mean and variance and which is trained with the negative log-likelihood; (iii) MC-DropoutNO (Gal & Ghahramani, 2016), where dropout in the feed-forward layers is used as approximate variational inference; and (iv) EnsembleNO (Lakshminarayanan et al., 2017), where we train $K$ randomly initialized FNO models (in our experiments, $K=10$) and compute the empirical mean and variance of the predictions.
@@ 5.1 Uncertainty estimates from DiverseNO
In Section 3, we showed that EnsembleNO outputs better uncertainty estimates OOD due to the diversity in the predictions among each model in the ensemble. Here, we evaluate whether DiverseNO can output good uncertainty estimates due to the diversity enforced over the last layer alone. Figure 4 shows the n-MeRCI metric for all UQ methods on increasing OOD shifts on the “easy”, ”medium” and “hard” cases of GPME. DiverseNO consistently outperforms BayesianNO, VarianceNO and MC-DropoutNO by 2×2$\times$2 × to 70×70$\times$70 × across all PDEs and OOD shifts, and is comparable (sometimes better) to the more expensive EnsembleNO that has diversity in weights of all layers (as seen in Figure 3). The n-MeRCI metric is generally close to zero for both methods indicating that their uncertainty estimates are well-correlated with the prediction errors and can be used to detect OOD shifts. See Section F.1 for additional metrics, solution profiles and uncertainty estimates.
@@ 5.2 Computational savings from DiverseNO
Figure 5: Cost-performance tradeoff. MSE normal-$\downarrow$ and n-MeRCI normal-$\downarrow$ vs. the number of floating point operations (FLOPs) for EnsembleNO and DiverseNO with varying number of parameters on the PME task with medium OOD shift, i.e., $m^{\text{train}}\in[2,3],m^{\text{test}}\in[4,5]$.
@@ Appendix A Related Work
Numerical methods. PDEs are ubiquitous throughout science and engineering, where they are used to model the evolution of various physical phenomena. These equations are typically solved for different values of PDE physical parameters, e.g., diffusivity in the heat equation, wavespeed in the advection equation, and the Reynolds number in Navier-Stokes equations. Solving PDEs typically requires extensive numerical knowledge and computational effort. Traditional approaches to solve PDEs (e.g., finite difference LeVeque (2007), finite volume LeVeque (2002) and finite element Hughes (2000) methods) can be computationally expensive, as their accuracy is dependent on the level of discretization of the spatial and temporal domains. Finer meshes are required to achieve high accuracy, resulting in increased computational costs. In addition, these approaches require a full re-run from scratch whenever there are changes in PDE parameters, which may not be known a priori.
@@
To alleviate the drawbacks of numerical methods, recent works in scientific machine learning (SciML) propose to use data-driven approaches to solve PDEs. These include so-called Physics-informed Neural Networks (PINNs) (Raissi et al., 2019), NOs (Li et al., 2020a, 2022b; Lu et al., 2019; Gupta et al., 2021; Yin et al., 2022), and reduced-order models for discovery (Brunton et al., 2016; Conti et al., 2023). By now, it has been shown that PINNs have several fundamental challenges associated with its soft constraint approach. In particular, it solves a single instance of the PDE with a fixed set of PDE parameters, is challenging to optimize for PDEs with large parameter values (Krishnapriyan et al., 2021; Edwards, 2022); and may return trivial solutions (Leiteritz & Pflüger, 2021). On the other hand, NOs (Kovachki et al., 2021; Li et al., 2020a) enjoy appealing properties of discretization invariance and universal approximation, while also achieving low approximation errors on in-domain tasks. Being purely data-driven, they are not guaranteed to satisfy all the physical properties of the solution. To address this, existing work has tried to incorporate different physical constraints via regularization (Li et al., 2021), within the architecture (e.g., boundary constraints in Saad et al. (2023), invariance in Liu et al. (2023), and PDE hard constraints in Négiar et al. (2023)) or via a projection to enforce conservation laws in Hansen et al. (2023). Most of these methods do not address the OOD problem that can occur even after enforcing these constraints. Subramanian et al. (2023) show that fine-tuning FNO models on OOD data is typically required to achieve reasonable performance. In particular, for significant OOD shifts, few-shot transfer learning requires a large amount of fine-tuning OOD data that may be unavailable for certain applications. Benitez et al. (2023) propose a variant of FNO specifically designed to learn the wavespeed to solution mapping in the Helmholtz equation and show that it performs better OOD.
@@
Several Bayesian deep learning methods common for standard neural networks have been shown to work well for NOs on in-domain PDE applications (Psaros et al., 2023; Zou et al., 2023). Commonly used methodologies for approximate Bayesian inference, e.g., the Bayesian Neural Operator (Magnani et al., 2022), DeepEnsembles (Lakshminarayanan et al., 2017), variational inference methods, e.g., Mean-field VI  (Graves, 2011; Teye et al., 2018) or MC-Dropout  (Gal & Ghahramani, 2016) and MCMC approaches, e.g., Hamiltonian Monte Carlo (HMC), provide, at best, crude approximations of the true posterior distribution in Equation 3 of this Bayesian model. For instance, DeepEnsembles train the same architecture multiple times to obtain different models that maximize the posterior, i.e., different modes of the posterior. The Bayesian Neural Operator uses a Laplace approximation of the posterior. Variational inference approaches approximate the posterior with a density $q({\mathbf{\theta}})$, and sample ${\mathbf{\theta}}_{j}\sim q({\mathbf{\theta}})$. MCMC methods construct a Markov chain that is asymptotically guaranteed to sample from the true posterior. While these methods can provide good in-domain uncertainty estimates, most of the scalable UQ methods are not robust to OOD shifts.
@@
There are several works that study the importance of diversity in ensemble-based approaches. Theisen et al. (2023) show that disagreement is key for an ensemble to be effective (with respect to accuracy). Wood et al. (2023) study the role of diversity in reducing in-domain generalization error. Diversity measures directly diversify the outputs from the different models in the ensemble, e.g., via minimizing the mutual information for classification tasks (Lee et al., 2022; Sinha et al., 2020; Rame & Cord, 2021), or the $L_{2}$ distance between the outputs for regression tasks. A recent work, DivDis (Lee et al., 2022) is the most related to our work in that it uses multiple prediction heads with diverse outputs to improve OOD accuracy in classification tasks. In particular, DivDis minimizes the mutual information between the outputs from different prediction heads when given OOD inputs (assumed to be known during training). This differs from our work since in operator learning, diversifying outputs directly does not provide informative uncertainty estimates.

### v2

@@ 1 Introduction
NOs have been shown to approximate the ground truth solution operator well for in-domain tasks  (Lu et al., 2019; Li et al., 2020a; Saad et al., 2023; Alesiani et al., 2022). However, to be useful, they also need to be robust for out-of-domain (OOD) applications. The robustness of OOD predictions is necessary in practical applications, as the test-time PDE parameters are typically not known during training. Figure 1 shows the sharp moving (discontinuous) shock solution to the “hard” Stefan problem, where the PDE parameter $u^{*}$ denotes the solution value at the shock point. We see that while a NO model trained to map the input PDE parameter $u^{*}$ to the solution of the Stefan problem accurately captures the sharp dynamics for in-domain values of $u^{*}$, it fails for OOD values of $u^{*}$, i.e., those values of $u^{*}$ that are outside the range covered by the training data. This failure shows that NOs are not robust in OOD scenarios and thus may not effectively capture the “true” underlying operator map in practical applications. It is important when developing these models to use reliable mechanisms to detect and correct such inaccurate predictions, as this limits their widespread deployment in critical scientific applications.
@@
Figure 1: 1-d Stefan Equation. In-domain (a) and OOD (b) predictions from FNO with post-hoc uncertainty estimates (3 standard deviations) obtained by BayesianNO (Magnani et al., 2022) and the proposed method at time $t=0.5$. BayesianNO outputs an accurate solution with reasonable uncertainty bounds in-domain (a), where the training and test input parameters are drawn from the same distribution, i.e., $u{{}^{*}}^{{\text{train}}},u{{}^{*}}^{{\text{test}}}\in[0.6,0.7]$. It fails to do so when the test inputs are OOD (b), with $u{{}^{*}}^{{\text{test}}}\in[0.5,0.55]$. The large oscillation occurs at different locations for different seeds. Proposed method (red) outputs accurate solution in-domain and improves the OOD predictions.
@@ 2.2 Sensitivity Analysis for Operator Learning
$p(u^{(i)}|{\mathbf{\theta}},\phi^{(i)})=\mathcal{N}(u^{(i)};{A}(\phi^{(i)};{\mathbf{\theta}}),\frac{1}{\alpha^{2}}\|{{A}^{\dagger}}(\phi^{(i)})\|^{2}_{2}{\mathbf{I}}).$
@@
${p({\mathbf{\theta}}|{\mathcal{D}}^{\text{train}})}\propto\prod_{i=1}^{N}p(u^{(i)}|{\mathbf{\theta}},\phi^{(i)})p({\mathbf{\theta}}).$ (2)
@@ 3.2 How do Ensembles Provide Better OOD UQ
The intermediate outputs from the different models in the ensemble also show diversity in operator learning until the last layer with a key difference between the in-domain and OOD final outputs. Given the same in-domain input, the first feed-forward (lifting) layer of these models already produces diverse outputs which continues until the penultimate layer (Figure 3(c)). Figure 3(d) shows that the last layer removes the diversity and maps the outputs to the same ground truth for an in-domain input. Figures 3(e) and 3(f) illustrates that a different trend occurs for an OOD input. The OOD intermediate outputs are also diverse, the difference here is that this diversity does not vanish from the last layer outputs since the model is not able to match the ground truth data. This representative example demonstrates how the ensemble model captures epistemic uncertainty in the OOD setting. We see that the individual models in the ensemble fit the training data with close to zero error, but are diverse and disagree for OOD inputs. Ensembles can approximate the BMA in Equation 3 with functionally diverse models from the posterior, leading to better OOD uncertainty estimates.
@@ 4.1 DiverseNO: Computationally Efficient OOD UQ
For the architecture change, we modify the last feed-forward layer to have $M$ output/prediction heads instead of one. (See Section C.1 for a hyperparameter study on the choice of $M$.) The architecture may be viewed as an ensemble with the individual models sharing parameters up to the penultimate layer. While parameter-sharing significantly reduces the computational complexity, when compared to using a full ensemble, it simultaneously hinders the diversity within the ensemble, which we showed is a crucial element for generating good uncertainty estimates (see Section 3.2).
To encourage diverse predictions, we propose to maximize the following diversity measure among the last-layer weights corresponding to the $M$ prediction heads. We also constrain the in-domain predictions from each of these heads to match the ground truth outputs. Formally, we solve:
@@
where ${\mathbf{\theta}}_{m}$ and ${\mathbf{\theta}}_{k}$ denote the last-layer weights corresponding to $m$-th and $k$-th prediction heads, respectively, $\hat{u}_{m}^{(i)}$ denotes the prediction from $m$-th output head for the $i$-th training example, and $u^{(i)}$ denotes the corresponding ground truth. The first term in Equation 4 is the relative $L_{2}$ loss standard in NO training. We add the regularizing second term to encourage diversity in the last-layer weights corresponding to the different head. The hyperparameter $\lambda_{\text{diverse}}$ controls the strength of the diversity regularization relative to the prediction loss. A naive selection procedure for $\lambda_{\text{diverse}}$ using only in-domain validation MSE selects an unconstrained model with no diversity penalty. Instead, we select the maximum regularization strength $\lambda_{\text{diverse}}$ that achieves an in-domain validation MSE within 10% of the identified best in-domain validation MSE. 10% is an arbitrary tolerance that denotes the % accuracy the practitioner is willing to forego to achieve better OOD UQ. This procedure trades off in-domain prediction errors for higher diversity that is useful for OOD UQ. See Section C.2 for a hyperparameter study on the diversity regularization strength $\lambda_{\text{diverse}}$; uncertainty metrics monotonically improve with higher $\lambda_{\text{diverse}}$. Section C.2 also includes ablations with regularizations for ensemble NNs that directly diversify the predictions (Bourel et al., 2020; Zhang et al., 2020).
@@ 4.2 Operator-ProbConserv
    Use the update rule from ProbConserv, described in Equation 6 below, to improve to the model.
@@ 5 Empirical Results
We use the Fourier Neural Operator (FNO) (Li et al., 2022a) as a base model and leave similar investigations of other operators to future work. To provide uncertainty estimates in PDE applications, we compare our proposed UQ method, DiverseNO  [^2] The code is available at https://github.com/amazon-science/operator-probconserv. with the following four commonly-used baselines of approximate Bayesian inference (Psaros et al., 2023): (i) BayesianNO (Magnani et al., 2022), that uses the (last-layer) Laplace approximation over the MAP estimate to approximate the posterior; (ii) VarianceNO (Lakshminarayanan et al., 2017), which outputs the mean and variance and which is trained with the negative log-likelihood; (iii) MC-DropoutNO (Gal & Ghahramani, 2016), where dropout in the feed-forward layers is used as approximate variational inference; and (iv) EnsembleNO (Lakshminarayanan et al., 2017), where we train $K$ randomly initialized FNO models (in our experiments, $K=10$) and compute the empirical mean and variance of the predictions.
@@ 5.1 Uncertainty estimates from DiverseNO
In Section 3, we showed that EnsembleNO outputs better uncertainty estimates OOD due to the diversity in the predictions among each model in the ensemble. Here, we evaluate whether DiverseNO can output good uncertainty estimates due to the diversity enforced over the last layer alone. Figure 4 shows the n-MeRCI metric for all UQ methods on increasing OOD shifts on the “easy”, ”medium” and “hard” cases of GPME. DiverseNO consistently outperforms BayesianNO, VarianceNO and MC-DropoutNO by 2×2$\times$2 × to 70×70$\times$70 × across all PDEs and OOD shifts, and is comparable (sometimes better) to the more expensive EnsembleNO that has diversity in weights of all layers (as seen in Figure 3). The n-MeRCI metric is generally close to zero for both methods indicating that their uncertainty estimates are well-correlated with the prediction errors and can be used to detect OOD shifts. See Section F.1 for additional metrics, solution profiles, uncertainty estimates and test problems, e.g., non-constant input and 2-d Darcy flow.
@@ 5.2 Computational savings from DiverseNO
Figure 5: Cost-performance tradeoff. MSE $\downarrow$ and n-MeRCI $\downarrow$ vs. the number of floating point operations (FLOPs) for EnsembleNO and DiverseNO with varying number of parameters on the PME task with medium OOD shift, i.e., $m^{\text{train}}\in[2,3],m^{\text{test}}\in[4,5]$.
@@ Appendix A Related Work
Numerical methods. PDEs are ubiquitous throughout science and engineering, where they are used to model the evolution of various physical phenomena. These equations are typically solved for different values of PDE physical parameters, e.g., diffusivity in the heat equation, wavespeed in the advection equation, and the Reynolds number in Navier-Stokes equations. Solving PDEs typically requires extensive numerical knowledge and computational effort. Traditional approaches to solve PDEs (e.g., finite difference (LeVeque, 2007), finite volume (LeVeque, 2002) and finite element (Hughes, 2000) methods) can be computationally expensive, as their accuracy is dependent on the level of discretization of the spatial and temporal domains. Finer meshes are required to achieve high accuracy, resulting in increased computational costs. In addition, these approaches require a full re-run from scratch whenever there are changes in PDE parameters, which may not be known a priori.
@@
To alleviate the drawbacks of numerical methods, recent works in scientific machine learning (SciML) propose to use data-driven approaches to solve PDEs. These include so-called Physics-informed Neural Networks (PINNs) (Raissi et al., 2019), NOs (Li et al., 2020a, 2022b; Lu et al., 2019; Gupta et al., 2021; Yin et al., 2022), and reduced-order models for discovery (Brunton et al., 2016; Conti et al., 2023). By now, it has been shown that PINNs have several fundamental challenges associated with its soft constraint approach. In particular, it solves a single instance of the PDE with a fixed set of PDE parameters; is challenging to optimize for PDEs with large parameter values (Krishnapriyan et al., 2021; Edwards, 2022); and may return trivial solutions (Leiteritz & Pflüger, 2021). On the other hand, NOs (Kovachki et al., 2021; Li et al., 2020a) enjoy appealing properties of discretization invariance and universal approximation, while also achieving low approximation errors on in-domain tasks. Being purely data-driven, they are not guaranteed to satisfy all the physical properties of the solution. To address this, existing work has tried to incorporate different physical constraints via regularization (Li et al., 2021), within the architecture (e.g., boundary constraints in Saad et al. (2023), invariance in Liu et al. (2023), and PDE hard constraints in Négiar et al. (2023)) or via a projection to enforce conservation laws in Hansen et al. (2023). Most of these methods do not address the OOD problem that can occur even after enforcing these constraints. Subramanian et al. (2023) show that fine-tuning FNO models on OOD data is typically required to achieve reasonable performance. In particular, for significant OOD shifts, few-shot transfer learning requires a large amount of fine-tuning OOD data that may be unavailable for certain applications. Benitez et al. (2023) propose a variant of FNO specifically designed to learn the wavespeed to solution mapping in the Helmholtz equation and show that it performs better OOD.
@@
Several Bayesian deep learning methods common for standard neural networks have been shown to work well for NOs on in-domain PDE applications (Psaros et al., 2023; Zou et al., 2023). Commonly used methodologies for approximate Bayesian inference, e.g., the Bayesian Neural Operator (Magnani et al., 2022), DeepEnsembles (Lakshminarayanan et al., 2017), variational inference methods, e.g., Mean-field VI (Graves, 2011; Teye et al., 2018) or MC-Dropout (Gal & Ghahramani, 2016) and MCMC approaches, e.g., Hamiltonian Monte Carlo, provide, at best, crude approximations of the true posterior distribution in Equation 3 of this Bayesian model. For instance, DeepEnsembles train the same architecture multiple times to obtain different models that maximize the posterior, i.e., different modes of the posterior. In applications to weather forecasting, FourCastNet (Pathak et al., 2022) generates ensembles by perturbing the initial condition with Gaussian noise. The Bayesian Neural Operator (Magnani et al., 2022) uses a last-layer Laplace approximation of the posterior. Weber et al. (2024) propose to use a Laplace approximation of the posterior via the last Fourier layer (instead of the last linear layer) to capture the global structure. Variational inference approaches approximate the posterior with a density $q({\mathbf{\theta}})$, and sample ${\mathbf{\theta}}_{j}\sim q({\mathbf{\theta}})$. MCMC methods construct a Markov chain that is asymptotically guaranteed to sample from the true posterior. An alternate approach in Guo et al. (2023) uses a latent space representation that is assumed to be aware of the confidence of the input data in relation to the region where the training data is located to provide point-wise uncertainty estimates. Recently, Ma et al. (2024) have applied distribution-free conformal prediction methods, which have been used for standard classification/regression tasks (Diquigiovanni et al., 2022; Angelopoulos et al., 2023), to NOs in the function space to provide uncertainty estimates for all points. While these methods can provide good in-domain uncertainty estimates, most of the scalable UQ methods are not robust to OOD shifts or have not been tested on OOD problems.
@@
There are several works that study the importance of diversity in ensemble-based approaches. Theisen et al. (2023) show that disagreement is key for an ensemble to be effective (with respect to accuracy). Wood et al. (2023) study the role of diversity in reducing in-domain generalization error. Diversity measures directly diversify the outputs from the different models in the ensemble, e.g., via minimizing the mutual information for classification tasks (Lee et al., 2022; Sinha et al., 2020; Rame & Cord, 2021), or the $L_{2}$ distance between the outputs for regression tasks. A recent work, DivDis (Lee et al., 2022) is the most related to our work in that it uses multiple prediction heads with diverse outputs to improve OOD accuracy in classification tasks. In particular, DivDis minimizes the mutual information between the outputs from different prediction heads when given OOD inputs (assumed to be known during training). This differs from our work since in operator learning, diversifying outputs directly does not provide informative uncertainty estimates. In applications to atmospheric forecasting, Lessig et al. (2023) propose AtmoRep, which is a Transformer-based foundation model that outputs multiple predictions by optimizing a loss function consisting of a MSE loss summed over the prediction heads, a Gaussian statistical loss and a variance regularization term that minimizes diversity across the prediction heads. In contrast, we show that low diversity can lead to poor uncertainty estimates on out-of-domain tasks and propose a regularization to maximize diversity in the weights of the prediction heads.